# Towards Wireless-Native Big AI Model: Insights into Its Ambitions, Peculiarities and Methodologies

### v1

@@ 1 Introduction
Given the outstanding potential of AI technology in complex feature extraction, high-dimensional data representation and adaptive decision-making, how to apply AI technology to solve wireless problems has been a critical topic in wireless technology research [1, 2, 3, 4]. However, simple migrating or reusing successful approaches in computer vision (CV) or natural language processing (NLP) to wireless systems tends to face serious bottlenecks in terms of performance and generalizability [5]. This “incompatibility” is due to the fact that AI for wireless is not directly equivalent to AI for natural language or computer vision. Specifically, AI for wireless essentially requires to learn electromagnetic physics and engineering methods, which makes it quite different from modeling human cognition and interaction capabilities in NLP and CV. Therefore, high-quality wireless AI technologies need to adapt to the characteristics of wireless data and the unique requirements of wireless systems, a concept known as ‘wireless-native’. Recent years, the incorporation of wireless-native methods such as physics-inspired structure [5, 6, 7, 8, 9, 10], and cross-module combination [11, 12] has significantly improved the performance and usability of wireless AI.
@@ 2.2 LLM-Based Wireless Agent
Although LLM-based methods have been somewhat explored, the current technical progress primarily focuses on instruction design, resource scheduling, and a few simple wireless problems. These approaches still lack effective solutions for numerous tough wireless challenges, such as scenario sensing and high-rate transmission. Therefore, instead of embracing LLMs, some researchers have explored BAIM in a wireless-native way. Wireless-native AI refers to an intelligent framework based on regular patterns of AI technologies, such as deep neural networks, while incorporating the peculiarities of wireless systems. It tailors model design to the nature of wireless data and system requirements, and adapts through re-learning on wireless data. Conventional wireless AI has also followed this approach in many studies. For example, in model design, researchers initially leveraged the numerical characteristics of wireless data, such as exploiting channel sparsity in the angle-delay domain and transforming it into the angle domain to make it more image-like [45, 46]. Later, researchers delved deeper into the physical mechanisms underlying the numerical characteristics and employed more fundamental, physics-inspired techniques to guide model optimization [5, 6, 7, 8, 10]. In terms of architecture, some cross-module [11, 12] and cross-layer [47] joint techniques have also been developed to facilitate applications in wireless systems.
@@ 3.2 Scientific Significance of wBAIM
Given these substantial differences, an in-depth understanding of electromagnetic laws is essential to more freely and accurately process, represent, and utilize electromagnetic signals. This necessitates the development of a BAIM that is native to electromagnetic sciences and systems—referred to as wBAIM. However, the black-box nature of AI models, coupled with the limitations of human cognition, makes it impossible to directly assess whether a given model has truly acquired the required capabilities. As a result, implicit assessment metrics become essential. This is precisely on of the reasons why [13] defines multi-task integration, multi-scenario unification, and all-in-one scheduling as the key features of wBAIM. If a single model can be used to generalize across multiple wireless tasks and scenarios, it must inherently capture the commonalities among them—that is, the general electromagnetic laws. In other words, this is the “compression is intelligence” in the wireless context.
@@ 4.3 Applied Scenarios
The unique networking architecture of wireless systems also brings peculiarities to wBAIM. On the one hand, since the major objective of wBAIM is to unify various scenarios and provide all-in-one scheduling, it requires wBAIM to be generalizable across diverse and heterogeneous cellular scenarios. The transmission configurations and scatterer distributions within different scenarios are usually different, making it a key technical challenge to enable wBAIM to quickly adapt to various scenarios. On the other hand, as wireless networks are evolving towards decentralization, the wireless data and computational resources are inherently distributed across scenarios. This decentralized nature makes it impractical for wBAIM to perform training and inference in a fully centralized manner, which differs from the mainstream LLM paradigm. Therefore, developing an efficient, multi-node architecture for collaborative computation is another critical technical requirement for advancing wBAIM.
@@ 5.1 Hybrid Data Collection
A rich knowledge base of training data is the foundation for BAIM to achieve wide-ranging intelligence. However, since wBAIM aims to serve massive heterogeneous cellular scenarios, the wireless datasets used for pre-training must adequately reflect the scenario diversity, which requires mixing the data from multiple scenarios. Meanwhile, to maximize the diversity of the dataset, an initial assessment of the similarity between scenarios, such as the data distribution sampling [51], should be performed before the data collection, and the scenarios with more significant differences should be prioritized.
@@ 5.2 Physics-Driven Learning Paradigm
To build broad-domain generalization intelligence for BAIM, it is essential to establish a learning paradigm that enables the model to learn various regular phenomena. In wireless networks, the regulations of different tasks and scenarios are derived from the physical laws of electromagnetic waves. Therefore, the ideal learning paradigm for wBAIM should be physics-driven. On the one hand, incorporating known physical laws into the model’s a priori design can significantly improve learning efficiency and generalization ability, reducing parameter redundancy and effectively controlling inference complexity. For instance, one could design the neural network’s representation structure according to the physical structure of wireless data, define the optimization objective based on the distribution of wireless features, or select the activation function based on the mapping relationship of wireless information.
@@ 5.4 Structural Prompt
After learning a wealth of knowledge and intelligence, the model’s efficient adaption to specific use cases is crucial for the deployment and application of BAIM. Common adaptation methods include fine-tuning, feature extraction, and prompting. Both fine-tuning and feature extraction require back-propagation and model retraining, which can be demanding on computational resources. Therefore, prompting, which does not require retraining, is the preferred adaptation method for wireless systems with highly decentralized computational resources.
@@ 5.5 Collaboration with AI-Agents
After constructing wBAIM, the next step is to consider the integration with AI-agent, an important booster for expanding the practical applications of current BAIM technology. Building upon wBAIM’s deep understanding of wireless regulations, adding modules such as planning, observation, memorization, and small model tool libraries will create wireless-native intelligent agent, significantly enhancing the usability and flexibility of wireless intelligence. For instance, a planning module can improve wBAIM’s logical reasoning capabilities and optimize its global performance; an observation module can enable autonomous acquisition of necessary state information, reducing signaling overhead; a memory module can help the agent to evolve continuously and reduce repetitive thinking; and small-tool libraries can lower the energy consumption and inference latency required for intelligent computations.
@@ 6 Conclusion
Acknowledgement  This work was supported in part by National Natural Science Foundation of China under Grants 62394292 and U20A20158, Ministry of Industry and Information Technology under Grant TC220H07E, Zhejiang Provincial Key R&D Program under Grant 2023C01021, and the Fundamental Research Funds for the Central Universities No. 226-2024-00069.

### v2

@@ 1 Introduction
Given the outstanding potential of AI technology in complex feature extraction, high-dimensional data representation and adaptive decision-making, how to apply AI technology to solve wireless problems has been a critical topic in wireless technology research [1, 2, 3, 4, 5, 6]. However, simple migrating or reusing successful approaches in computer vision (CV) or natural language processing (NLP) to wireless systems tends to face serious bottlenecks in terms of performance and generalizability [7]. This “incompatibility” is due to the fact that AI for wireless is not directly equivalent to AI for natural language or computer vision. Specifically, AI for wireless essentially requires to learn electromagnetic physics and engineering methods, which makes it quite different from modeling human cognition and interaction capabilities in NLP and CV. Therefore, high-quality wireless AI technologies need to adapt to the characteristics of wireless data and the unique requirements of wireless systems, a concept known as ‘wireless-native’. Recent years, the incorporation of wireless-native methods such as physics-inspired structure [7, 8, 9, 10, 11, 12], and cross-module combination [13, 14] has significantly improved the performance and generalization of wireless AI. Meanwhile, the introduction of prior information and advanced learning architectures can also reduce the demand for training data and computational resources. These fundamental advantages are of great significance for advancing the usability of wireless AI, whether in centralized or distributed usage scenarios [2].
@@ 2.2 LLM-Based Wireless Agent
Although LLM-based methods have been somewhat explored, the current technical progress primarily focuses on instruction design, resource scheduling, and a few simple wireless problems. These approaches still lack effective solutions for numerous tough wireless challenges, such as scenario sensing and high-rate transmission. This limitation arises from two primary factors: firstly, linguistic expressions are frequently insufficient in precisely characterizing multimodal signals and their intricate correlations; secondly, the language intelligence derived from textual training lacks adequate knowledge reserves to comprehend the data structures and underlying mechanisms inherent in wireless data. Therefore, instead of embracing LLMs, some researchers have explored BAIM in a wireless-native way. Wireless-native AI refers to an intelligent framework based on regular patterns of AI technologies, such as deep neural networks, while incorporating the peculiarities of wireless systems. It tailors model design to the nature of wireless data and system requirements, and adapts through re-learning on wireless data. Conventional wireless AI has also followed this approach in many studies. For example, in model design, researchers initially leveraged the numerical characteristics of wireless data, such as exploiting channel sparsity in the angle-delay domain and transforming it into the angle domain to make it more image-like [47, 48]. Later, researchers delved deeper into the physical mechanisms underlying the numerical characteristics and employed more fundamental, physics-inspired techniques to guide model optimization [7, 8, 9, 10, 12]. In terms of architecture, some cross-module [13, 14] and cross-layer [49] joint techniques have also been developed to facilitate applications in wireless systems.
@@ 3.2 Scientific Significance of wBAIM
Given these substantial differences, an in-depth understanding of electromagnetic laws is essential to more freely and accurately process, represent, and utilize electromagnetic signals. This necessitates the development of a BAIM that is native to electromagnetic sciences and systems—referred to as wBAIM. However, the black-box nature of AI models, coupled with the limitations of human cognition, makes it impossible to directly assess whether a given model has truly acquired the required capabilities. As a result, implicit assessment metrics become essential. This is precisely on of the reasons why [15] defines multi-task integration, multi-scenario unification, and all-in-one scheduling as the key features of wBAIM. If a single model can be used to generalize across multiple wireless tasks and scenarios, it must inherently capture the commonalities among them—that is, the general electromagnetic laws. In other words, this is the “compression is intelligence” in the wireless context. Likely, this intelligence not only serves wireless communication systems but also provides technical supports for other electromagnetics-related fields such as remote sensing [54, 55] and radar systems [56].
@@ 4.3 Applied Scenarios
The unique networking architecture of wireless systems also brings peculiarities to wBAIM. On the one hand, since the major objective of wBAIM is to unify various scenarios and provide all-in-one scheduling, it requires wBAIM to be generalizable across diverse and heterogeneous cellular scenarios. The transmission configurations and scatterer environments within different scenarios are usually different, making it a key technical challenge to enable wBAIM to quickly adapt to various scenarios. On the other hand, as wireless networks are evolving towards decentralization, the wireless data and computational resources are inherently distributed across scenarios. This decentralized nature makes it impractical for wBAIM to perform training and inference in a fully centralized manner, which differs from the mainstream LLM paradigm. Therefore, developing an efficient, multi-node architecture for collaborative computation is another critical technical requirement for advancing wBAIM.
@@ 5.1 Hybrid Data Collection
A rich knowledge base of training data is the foundation for BAIM to achieve wide-ranging intelligence. However, since wBAIM aims to serve massive heterogeneous cellular scenarios, the wireless datasets used for pre-training must adequately reflect the scenario diversity, which requires mixing the data from multiple scenarios. Meanwhile, to maximize the diversity of the dataset, an initial assessment of the similarity between scenarios, such as the data distribution sampling [53], should be performed before the data collection, and the scenarios with more significant differences should be prioritized. In addition, given privacy and security considerations, federated learning [57], which employs model interaction rather than direct data mixing, represents a potential technical approach.
@@ 5.2 Physics-Driven Learning Paradigm
To build broad-domain generalization intelligence for BAIM, it is essential to establish a learning paradigm that enables the model to learn various regular phenomena. In wireless networks, the regulations of different tasks and scenarios are derived from the physical laws of electromagnetic waves. Therefore, the ideal learning paradigm for wBAIM should be physics-driven. On the one hand, incorporating known physical laws into the model’s a priori design can significantly improve learning efficiency and generalization ability, reducing parameter redundancy and effectively controlling inference complexity. For instance, one could design the neural network’s representation structure according to the physical structure of wireless data, define the optimization objective based on the distribution of wireless features, or select the activation function based on the mapping relationship of wireless information [8].
@@ 5.4 Structural Prompt
After learning a wealth of knowledge and intelligence, the model’s efficient adaption to specific use cases is crucial for the deployment and application of BAIM. Common adaptation methods include fine-tuning, feature extraction, and prompting. Both fine-tuning and feature extraction require back-propagation and gradient descent in batches, which are demanding on hardware devices and computational resources. Meanwhile, this retraining process often demands considerable time, during which the model’s intelligence expression may fail to meet current scenario requirements - a particularly critical issue for wireless communication systems with frequent environmental and configuration changes. Therefore, prompting, which does not require retraining, is the preferred adaptation method for wireless systems with highly decentralized computational resources.
@@ 5.5 Collaboration with AI-Agents
After constructing wBAIM, the next step is to consider the integration with AI-agent, an important booster for expanding the practical applications of current BAIM technology. Building upon wBAIM’s deep understanding of wireless regulations, adding modules such as planning, observation, memorization, and small model tool libraries will create wireless-native intelligent agent, significantly enhancing the usability and flexibility of wireless intelligence. For instance, a planning module can improve wBAIM’s logical reasoning capabilities and optimize its global performance; an observation module can enable autonomous acquisition of necessary state information, reducing signaling overhead; a memory module can help the agent to evolve continuously and reduce repetitive thinking; and small-tool libraries can lower the energy consumption required for intelligent computations and optimize the system configurations. An overall example is shown in Fig. 7. Furthermore, the technical crux of coordinating such multifunctional modules lies in controlling the overall computational latency and achieving effective fusion and alignment of multi-modal information.
Figure 7: An overall example of wBAIM’s collaboration with AI-Agents.
@@ 6 Conclusion
Acknowledgement  This work was supported in part by National Natural Science Foundation of China under Grants 62394292, U20A20158 and 624B2129, Ministry of Industry and Information Technology under Grant TC220H07E, Zhejiang Provincial Key R&D Program under Grant 2023C01021, and the Fundamental Research Funds for the Central Universities No. 226-2024-00069.